# Imprecise Markov Semigroups and their Ergodicity

### v4

@@
Abstract.
We introduce the concept of an imprecise Markov semigroup $\mathbf{Q}$. It is a tool that allows us to represent ambiguity around both the initial and the transition probabilities of a continuous-time Markov process via a compact collection of Markov semigroups, each associated with a (possibly different) Markov process. We use techniques from topology, geometry, and probability to study the ergodic behavior of $\mathbf{Q}$. We show that, under some conditions that also involve the geometry of the state space, eventually the ambiguity fades. We call this property ergodicity of the imprecise Markov semigroup, and we relate it to the classical notion of ergodicity. We prove ergodicity both when the state space is Euclidean or a Riemannian manifold, and when it is an arbitrary measurable space. The importance of our findings for the fields of artificial intelligence and computer vision is also discussed, in particular in the study of how the probability of an output evolves over time as we perturb the input of a convolutional autoencoder.
Key words and phrases:
@@ Abstract.
1. Introduction
When dealing with a stochastic phenomenon whose future evolution is independent of its history, it is customary to model its behavior via a stochastic process enjoying the Markov property. Loosely, this means that the law of the realization $X_{T}$ of the process $(X_{t})_{t\geq 0}$ at time $T$, given what happened up to time $s<T$, is equivalent to the law of $X_{T}$, given only $X_{s}$. We can of course express this property in a slightly more general way, by saying that the expectation of a generic bounded measurable real-valued function $\tilde{f}$ of $X_{T}$, given what happened in the past up to time $s<T$, is equal to the expectation of $\tilde{f}(X_{T})$, given only $X_{s}$. [^1] Throughout the paper, we do not use the pronoun “we” as a form of pluralis majestatis, but rather to comply to the convention of scientific papers of writing in plural form. Notice that in the present paper we perpetrate a terminology abuse, and refer to real-valued functions on the state space $E$ as “functionals”. In addition, we only consider continuous-time processes $(X_{t})_{t\geq 0}$.
@@ 1. Introduction
Imprecise probability (Augustin et al., 2014; Troffaes and de Cooman, 2014; Walley, 1991) is a field whose techniques naturally lend themselves to be used for this purpose. In fact, imprecise Markov processes are a flourishing area of the imprecise probability literature (Crossman and Škulj, 2010; De Bock et al., 2021; de Cooman et al., 2009; Delgado et al., 2011; Jaeger et al., 2020; Krak et al., 2017, 2019; T’Joens and De Bock, 2021; Trevizan et al., 2007, 2008; Troffaes et al., 2019), [^2] Much of the existing literature focuses on discrete-time processes. with numerous applications e.g. in engineering (Aslett et al., 2022, Chapter 4), including aerospace engineering (Vasile, 2021, Chapter 5), and economics (Cerreia-Vioglio et al., 2023; Denk et al., 2018, 2020). As the name suggests, imprecise Markov processes allow to take into account imprecision in the initial law and the transition probabilities of a Markov process.
@@
In this paper, we introduce a new concept that we call an Imprecise Markov Semigroup. It is a compact collection of (precise) Markov semigroups, each associated with a Markov process. [^3] Here, by “precise” we refer to the classical notion of a Markov semigroup. Modeling the aforementioned types of imprecision via such a collection, instead of using lower previsions (Troffaes and de Cooman, 2014) or Choquet integrals (Choquet, 1954), allows us to develop new geometric and probabilistic techniques based on Bakry et al. (2014) that make the study of the limiting behavior of imprecise Markov processes (captured by the limit behavior of the imprecise Markov semigroup) easier. [^4] Since we consider a collection of precise Markov semigroups, one could say ours is a “precise approach” to imprecision. In particular, we discover a relationship between such a behavior and the curvature of the state space – that is, the space $E$ that the elements of the process $(X_{t})_{t\geq 0}$ take values on. This also relates the present work to the study of the geometry of uncertainty (Cuzzolin, 2021).
Our goal is to study what we call the $\tilde{f}-$ergodic property of an imprecise Markov semigroup. Roughly, we want to see whether a form of the classical ergodic mantra “in the limit, time average (of the form $\lim_{t\rightarrow\infty}\mathbb{E}[\tilde{f}(X_{t})\mid X_{0}]$) equals space average (of the form $\mathbb{E}[\tilde{f}(X_{0})]$)” can be derived (Cerreia-Vioglio et al., 2015; Caprio and Mukherjee, 2023), for some bounded functional $\tilde{f}$ of interest. Slightly more precisely, we inspect whether the limits as $t\rightarrow\infty$ of the maximum and the minimum values for the expectation $\mathbb{E}[\tilde{f}(X_{t})\mid X_{0}]$ – captured by the maximum and minimum elements of the imprecise Markov semigroup – are equal to the expectation $\mathbb{E}[\tilde{f}(X_{0})]$, taken with respect to some measures of interest. The answer, which – as we shall see – follows from Corollary 8.1 (where the state space $E$ is a Euclidean space or a Riemannian manifold) and Corollary 10.1 (where $E$ is an arbitrary measurable space), is positive, under some conditions (including some geometric ones, as we briefly mentioned earlier).
In addition, we discover that – under a further condition interpretable as the initial distribution of the process, i.e. the distribution of $X_{0}$, being known and invariant – in the limit the imprecision vanishes. We find that the limits for both the maximum and the minimum of $\mathbb{E}[\tilde{f}(X_{t})\mid X_{0}]$ coincide, and are equal to $\mathbb{E}[\tilde{f}(X_{0})]$ (taken with respect to the initial distribution of the process). The similarities and differences between our results and the existing ones in the imprecise Markov literature, and in imprecise ergodic theory (Caprio and Gong, 2023; Caprio and Mukherjee, 2023; Cerreia-Vioglio et al., 2015), are discussed in Sections 2 and 3.
@@
Besides being interesting in their own right (and because they provide new techniques to the fields of imprecise Markov processes), we believe that our findings will prove useful in future research in machine learning, artificial intelligence, and computer vision. Indeed, in recent works on convolutional autoencoders, images (or, more in general, inputs from an input space $\mathcal{X}$) are “projected” onto a low-dimensional manifold $E$ by an encoding function $\phi_{\text{enc}}:\mathcal{X}\rightarrow E$. Different “portions” of the surface $\mathcal{S}(E)$ of such a manifold capture different features of the input image (Yu et al., 2023). For instance, a portion of the surface may correspond to the feature “dog”, so that dog images are projected there. This means that we can consider a partition $\mathscr{E}=\{E_{j}\}_{j=1}^{J}$ of $\mathcal{S}(E)$ whose elements correspond to “feature portions”. For example, $J$ could be equal to the cardinality $|\mathcal{Y}|$ of the output space $\mathcal{Y}$ in the case of classification problems. Then, a decoding function $\phi_{\text{dec}}:E\rightarrow\text{Prob}(\mathcal{Y})$ returns the probabilities of $y_{j}\in\mathcal{Y}$ being the correct output for the input $x\in\mathcal{X}$. Here $\text{Prob}(\mathcal{Y})$ denotes the space of countably additive probabilities on $\mathcal{Y}$. For example, in a classification case, $(\phi_{\text{dec}}\circ\phi_{\text{enc}})(x)=(P(y_{1}),\ldots,P(y_{J}))^{\top}$, where $\mathcal{Y}=\{y_{1},\ldots,y_{J}\}$ is the label space, and $P(y_{j})$ is the probability estimated by the model that $y_{j}$ is the correct label for input $x$, $j\in\{1,\ldots,J\}$. A visual representation is given in Figure 1.
@@
Figure 1. In this figure, we provide a visual representation of the functioning of a convolutional autoencoder in an image classification problem. The input image of a pug is mapped by encoding function $\phi_{\text{enc}}$ to the surface area $E_{\text{dog}}\in\mathscr{E}$ corresponding to the feature “dog”. Such a projection is then mapped by decoding function $\phi_{\text{dec}}$ to the probability space over the labels. In this simple example the model gives a high probability to the first label, i.e. “dog”.
@@
Now, imagine that we want to study a smooth transition between a dog and a cat image. Intuitively, this corresponds to a “walk” from the “dog portion” $E_{\text{dog}}$ of the manifold’s surface $\mathcal{S}(E)$ to the “cat portion” $E_{\text{cat}}$ (Pérez, 1998). A visual example is given in Figure 2. Our results, and especially Corollary 8.1, may be very useful when (i) such a “walk” is not deterministic, but rather modeled via a Markov process; (ii) the user is uncertain about which initial and transition probabilities to choose; and (iii) we are interested in the probability of a particular label during such a walk. For example, if we are interested in how the probability of the first label evolves over time, we could consider function $z\mapsto\tilde{f}(z)=e_{1}^{\top}\phi_{\text{dec}}(z)$, where $e_{1}=(1,0,0,\ldots,0)^{\top}$ is a $J$-dimensional standard basis vector, and $\tilde{f}$ is an example of a bounded functional on $E$, as discussed earlier in this section.
@@
Figure 2. In this figure, we represent the smooth transition between a dog and a cat pictures. Moving a cursor between the pictures causes a walk from the “dog portion” $E_{\text{dog}}$ of the manifold’s surface $\mathcal{S}(E)$ to the “cat portion” $E_{\text{cat}}$, depicted as a purple smooth curve. The results we present in this paper are useful to determine the behavior of a similar walk that is not deterministic, when it is not possible to determine unique initial and transition probabilities, and when we are interested e.g. in the evolution of the probability of the first label, $y_{1}=\text{``dog''}$, captured by the functional $z\mapsto\tilde{f}(z)=e_{1}^{\top}\phi_{\text{dec}}(z)$.
@@
Our paper is structured as follows. Section 2 introduces the concept of an Imprecise Markov Semigroup. Section 3 derives our results when state space $E$ is a Euclidean space or a Riemannian manifold. Section 4 generalizes the findings in Section 3 to the case of $E$ being an arbitrary measurable space. Section 5 concludes our work. We prove Theorem 10 in A, and we provide a primer on Markov semigroups and diffusion Markov triples in B.
2. Imprecise Markov Semigroups
@@
In this section, we introduce the concept of an Imprecise Markov Semigroup. Its definition hinges on the notions of Markov processes, Markov semigroups, carré du champ and iterated carré du champ operators, and Bakry-Émery curvature. We refer the unfamiliar reader to B.1.
@@
$P_{t}f(x)=\mathbb{E}\left[f(X_{t})\mid X_{0}=x\right]=\int_{E}f(y)p_{t}(x,\text{d}y),\quad f\in B(E)\text{, }x\in E,$
where $B(E)$ is the space of bounded measurable functionals on $E$, and $p_{t}$ is the probability kernel describing the evolution of the Markov process. In words, operator $P_{t}$ applied on $f\in B(E)$ gives us the expected value of $f(X_{t})$, “weighted by” the transition probability (from the starting value $X_{0}=x$ to $X_{t}$) of our Markov process. We focus on this type of Markov semigroups because studying their (ergodic) behavior inform us about the (ergodic) behavior of the associated Markov processes.
Throughout the paper, we will assume that the semigroups we consider always have an invariant probability measure $\mu$, so that the elements of $\mathbf{P}=(P_{t})_{t\geq 0}$ are bounded operators in $\mathbb{L}^{p}(\mu)$, $p\in[1,\infty)$. In fact, most semigroups of interest have an invariant measure (Bakry et al., 2014, Section 1.2.1). A way to derive $\mu$ as a weak limit of a “reasonable” initial probability measure $\mu_{0}$ is inspected in Bakry et al. (2014, Page 10). For a discussion on this matter, we refer the interested reader to Remark 2 in B.1.
@@
Consider now a collection $\mathbf{Q}$ of Markov semigroups (MSGs) associated with Markov processes. That is, let $\mathbf{Q}\subset({B(E)}^{\mathbb{R}_{+}})^{B(E)}$ be such that
@@
The notion of Imprecise Markov Semigroup hinges on the possibility of comparing the elements of $\mathbf{Q}$, when they are evaluated at some functional $\tilde{f}\in B(E)$. First, suppose we are able to define a total order $\preceq^{\text{tot}}_{\tilde{f}}$ so that
@@
Being $\preceq^{\text{tot}}_{\tilde{f}}$ a total order, this means that, for all $\mathbf{P},\mathbf{P^{\prime}}\in\mathbf{Q}$, either $\mathbf{P}\preceq^{\text{tot}}_{\tilde{f}}\mathbf{P^{\prime}}$ or $\mathbf{P^{\prime}}\preceq^{\text{tot}}_{\tilde{f}}\mathbf{P}$. In addition, the notions of maximal and greatest elements of $(\mathbf{Q},\preceq^{\text{tot}}_{\tilde{f}})$ coincide, and correspond to the classical notion of maximum (with respect to the total order $\preceq^{\text{tot}}_{\tilde{f}}$). Similarly, minimal and least element notions coincide, and correspond to the classical notion of minimum.
@@
Now, let $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$ be the order topology. That is (provided $|\mathbf{Q}|\geq 2$), $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$ is the topology whose base is given by the open rays $\{\mathbf{P}:\mathbf{P}_{1}\prec^{\text{tot}}_{\tilde{f}}\mathbf{P}\}$ and $\{\mathbf{P}:\mathbf{P}\prec^{\text{tot}}_{\tilde{f}}\mathbf{P}_{2}\}$, for all $\mathbf{P}_{1}$, $\mathbf{P}_{2}$ in $\mathbf{Q}$, and by the open intervals $(\mathbf{P}_{1},\mathbf{P}_{2})=\{\mathbf{P}:\mathbf{P}_{1}\prec^{\text{tot}}_{\tilde{f}}\mathbf{P}\prec^{\text{tot}}_{\tilde{f}}\mathbf{P}_{2}\}$. This means that the $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$-open sets in $\mathbf{Q}$ are the sets that are a union of (possibly infinitely many) such open intervals and rays.
@@
Assume that $\mathbf{Q}$ is $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$-compact. This implies that $\mathbf{Q}$ contains the minimum and the maximum (with respect to the total order $\preceq^{\text{tot}}_{\tilde{f}}$), which we denote by $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$, respectively. In turn, since $\underline{\mathbf{P}},\overline{\mathbf{P}}\in\mathbf{Q}$, they both are (precise) Markov semigroups associated with Markov processes by (1).
@@
Suppose now that – contrary to before – we are only able to define a partial order $\preceq^{\text{part}}_{\tilde{f}}$ on $\mathbf{Q}$ as the previous one. More formally, we have that

### v5

# arXiv:2405.00081v5[math.PR] 02 Mar 2026

@@
Abstract
We introduce the concept of an imprecise Markov semigroup $\mathbf{Q}$. It is a tool that allows us to represent ambiguity around both the transition probabilities and the invariant measure of a continuous-time Markov process via a collection of Markov semigroups, each associated with a (possibly different) Markov process. We use techniques from topology, geometry, and probability to analyze ergodic limits under model uncertainty encoded by $\mathbf{Q}$. We establish long-term bounds that are uniform in the initial state and identify regimes in which the imprecision in these bounds collapses asymptotically. Our results are proved in progressively more general settings. We first assume that $\mathbf{Q}$ is compact and that the state space is Euclidean or a Riemannian manifold, working with a fixed bounded observable. We then allow the state space to be standard Borel, while keeping $\mathbf{Q}$ compact and the observable fixed. Finally, we drop compactness and work on Polish metric spaces of finite diameter, where we treat arbitrary bounded Lipschitz observables. The importance of our findings for the fields of artificial intelligence and computer vision is also discussed at a high level; In particular, in the study of how the probability of an output evolves over time as we perturb the input of a convolutional autoencoder.
keywords:
@@ Abstract
1 Introduction
When dealing with a stochastic phenomenon whose future evolution is independent of its history, it is customary to model its behavior via a stochastic process enjoying the Markov property. Loosely, this means that the law of the realization $X_{T}$ of the process $(X_{t})_{t\geq 0}$ at time $T$, given what happened up to time $s<T$, is equivalent to the law of $X_{T}$, given only $X_{s}$. We can of course express this property in a slightly more general way, by saying that the expectation of a generic bounded measurable real-valued function $\tilde{f}$ of $X_{T}$, given what happened in the past up to time $s<T$, is equal to the expectation of $\tilde{f}(X_{T})$, given only $X_{s}$. Notice that in the present paper we perpetrate a terminology abuse, and refer to real-valued functions on the state space $E$ as “functionals”. In addition, we only consider continuous-time processes $(X_{t})_{t\geq 0}$.
@@ 1 Introduction
Imprecise probability (Augustin et al., 2014; Troffaes and de Cooman, 2014; Walley, 1991) is a field whose techniques naturally lend themselves to be used for this purpose. In fact, imprecise Markov processes are a flourishing area of the imprecise probability literature (Crossman and Škulj, 2010; De Bock et al., 2021; de Cooman et al., 2009; Delgado et al., 2011; Jaeger et al., 2020; Krak et al., 2017, 2019; T’Joens and De Bock, 2021; Trevizan et al., 2007, 2008; Troffaes et al., 2019), [^1] Much of the existing literature focuses on discrete-time processes. with numerous applications e.g. in engineering (Aslett et al., 2022, Chapter 4), including aerospace engineering (Vasile, 2021, Chapter 5), and economics (Cerreia-Vioglio et al., 2023; Denk et al., 2018, 2020). As the name suggests, imprecise Markov processes allow to take into account imprecision in the initial and transition probabilities (and also in the invariant measure, when it exists) of a Markov process.
@@
In this paper, we introduce a new concept that we call an Imprecise Markov Semigroup. It is a collection $\mathbf{Q}$ of (precise) Markov semigroups, each associated with a Markov process. [^2] Here, by “precise” we refer to the classical notion of a Markov semigroup. Modeling the imprecision on the transition probabilities and on the invariant measure (whenever it exists) via such a collection, instead of using lower previsions (Troffaes and de Cooman, 2014) or Choquet integrals (Choquet, 1954), allows us to combine geometric and probabilistic techniques based on the Markov diffusion operators literature (Bakry et al., 2014) that make the study of the limiting behavior of imprecise Markov processes (captured by the limit behavior of the imprecise Markov semigroup) easier. In particular, we are able to relate such a behavior to the curvature of the state space – that is, the space $E$ that the elements of the process $(X_{t})_{t\geq 0}$ take values on. This also links the present work to the study of the geometry of uncertainty (Cuzzolin, 2021).
Our goal is to study what we call the ergodic property of an imprecise Markov semigroup. Throughout the paper we use “ergodicity” in the semigroup/long-term sense: We study the asymptotic behavior, as $t\to\infty$, of conditional expectations of observables under the Markov evolution, i.e. $P_{t}f(x)=\mathbb{E}[f(X_{t})\mid X_{0}=x]$. In Sections 3 and 4, the bounds we derive are formulated for a fixed (bounded) test functional $\tilde{f}$; Accordingly, the conclusions should be read as long-term bounds for the expectation of $\tilde{f}$ whose limiting constants do not depend on the initial state, rather than as full distributional convergence – which would correspond to convergence for all $f\in B(E)$. In contrast, in Section 5 we work on the larger class $\mathrm{Lip}_{b}(E)$ of bounded Lipschitz functionals and obtain uniform-in-$x$ limits for the static upper and lower envelopes $\overline{T}_{t}$ and $\underline{T}_{t}$ of $\mathbf{Q}$, respectively, and for the time-consistent (upper/lower) Nisio semigroups when switching is allowed. Such results are derived for every bounded Lipschitz observable $f$, which is closer in spirit to distributional stability. Finally, we emphasize that our limits are not Cesàro time averages: We do not prove statements of the form $\frac{1}{T}\int_{0}^{T}f(X_{t})\mathrm{d}t\to\int f\mathrm{d}\mu$. Rather, we establish convergence of the semigroup as $t\to\infty$ (equivalently, convergence of one-time marginals and/or of their upper/lower envelopes when switching is allowed). When appropriate (e.g. under standard invariance/ergodicity assumptions), classical ergodic theorems can then be invoked to relate such semigroup convergence to time average statements, but this is not the object of the present results.
The similarities and differences between our results and the existing ones in the imprecise Markov literature and in imprecise ergodic theory (Kozine and Utkin, 2002; Hermans and de Cooman, 2012; Skulj, 2013, 2015; Cerreia-Vioglio et al., 2015; de Cooman et al., 2016; Caprio and Gong, 2023; Caprio and Mukherjee, 2023), are discussed in Sections 2 and 3, and 5.
@@
Besides being interesting in their own right (and because they provide new techniques to the fields of imprecise Markov processes), we believe that our findings will prove useful in future research in machine learning, artificial intelligence, and computer vision. Let us add a high-level discussion on this topic. In recent works on convolutional autoencoders, images (or, more in general, inputs from an input space $\mathcal{X}$) are “projected” onto a low-dimensional manifold $E$ by an encoding function $\phi_{\text{enc}}:\mathcal{X}\rightarrow E$. Different “portions” of the surface $\mathcal{S}(E)$ of such a manifold capture different features of the input image (Yu et al., 2023). For instance, a portion of the surface may correspond to the feature “dog”, so that dog images are projected there. This means that we can consider a partition $\mathscr{E}=\{E_{j}\}_{j=1}^{J}$ of $\mathcal{S}(E)$ whose elements correspond to “feature portions”. For example, $J$ could be equal to the cardinality $|\mathcal{Y}|$ of the output space $\mathcal{Y}$ in the case of classification problems. Then, a decoding function $\phi_{\text{dec}}:E\rightarrow\Delta_{\mathcal{Y}}$ returns the probabilities of $y_{j}\in\mathcal{Y}$ being the correct output for the input $x\in\mathcal{X}$. Here $\Delta_{\mathcal{Y}}$ denotes the space of countably additive probabilities on $\mathcal{Y}$. For example, in a classification case, $(\phi_{\text{dec}}\circ\phi_{\text{enc}})(x)=(P(y_{1}),\ldots,P(y_{J}))^{\top}$, where $\mathcal{Y}=\{y_{1},\ldots,y_{J}\}$ is the label space, and $P(y_{j})$ is the probability estimated by the model that $y_{j}$ is the correct label for input $x$, $j\in\{1,\ldots,J\}$. A visual representation is given in Figure 1.
@@
Figure 1: In this figure, we provide a visual representation of the functioning of a convolutional autoencoder in an image classification problem. The input image of a pug is mapped by encoding function $\phi_{\text{enc}}$ to the surface area $E_{\text{dog}}\in\mathscr{E}$ corresponding to the feature “dog”. Such a projection is then mapped by decoding function $\phi_{\text{dec}}$ to the probability space over the labels. In this simple example the model gives a high probability to the first label, i.e. “dog”.
@@
Now, imagine that we want to study a smooth transition between a dog and a cat image. Intuitively, this corresponds to a “walk” from the “dog portion” $E_{\text{dog}}$ of the manifold’s surface $\mathcal{S}(E)$ to the “cat portion” $E_{\text{cat}}$ (Pérez, 1998). A visual example is given in Figure 2. Our results, and especially Corollaries 10.1 and 19.1, may be very useful when (i) such a “walk” is not deterministic, but rather modeled via a Markov process; (ii) the user is uncertain about which transition probabilities to choose (and on the nature of the invariant measure); and (iii) we are interested in the dynamics of the probability of a particular label during such a walk. For example, if we are interested in how the probability of the first label evolves over time, we could consider function $z\mapsto\tilde{f}(z)=e_{1}^{\top}\phi_{\text{dec}}(z)$, where $e_{1}=(1,0,0,\ldots,0)^{\top}$ is a $J$-dimensional standard basis vector, and $\tilde{f}$ is an example of a bounded functional on $E$, as discussed earlier in this section.
@@
Figure 2: In this figure, we represent the smooth transition between a dog and a cat pictures. Moving a cursor between the pictures causes a walk from the “dog portion” $E_{\text{dog}}$ of the manifold’s surface $\mathcal{S}(E)$ to the “cat portion” $E_{\text{cat}}$, depicted as a purple smooth curve. The results we present in this paper are useful to determine the behavior of a similar walk that is not deterministic, when it is not possible to determine exactly transition probabilities (and the invariant measure, when it exists), and when we are interested e.g. in the evolution of the probability of the first label, $y_{1}=\text{``dog''}$, captured by the functional $z\mapsto\tilde{f}(z)=e_{1}^{\top}\phi_{\text{dec}}(z)$.
@@
Our paper is structured as follows. Section 2 introduces the concept of an Imprecise Markov Semigroup $\mathbf{Q}$. Section 3 derives our results when state space $E$ is a Euclidean space or a Riemannian manifold, $\mathbf{Q}$ is compact, and the bounded test functional $\tilde{f}$ is fixed. Section 4 generalizes the findings in Section 3 to the case of $E$ being an arbitrary measurable space, while keeping $\mathbf{Q}$ compact and the bounded test functional $\tilde{f}$ fixed. Section 5 further extends our results: It is only required that $E$ is Polish metric with finite diameter and that the test map is any bounded Lipschitz functional. Section 6 concludes our work. We prove Theorem 12 in Appendix A, and we provide a primer on Markov semigroups and diffusion Markov triples in Appendix C.
2 Imprecise Markov Semigroups
@@
In this section, we introduce the concept of an Imprecise Markov Semigroup. Its definition hinges on the notions of Markov processes, Markov semigroups, carré du champ and iterated carré du champ operators, and Bakry-Émery curvature. We refer the unfamiliar reader to Appendix C.1.
@@
Here, the measure decomposition theorem refers to the disintegration theorem (existence of regular conditional distributions). Concretely, if $\mu$ is a probability measure on $(E\times E,\mathcal{F}\otimes\mathcal{F})$ and $\mu_{1}$ is its first marginal, then there exists a Markov kernel $k(x,\text{d}y)$ (measurable in $x$) such that $\mu(\text{d}x,\text{d}y)=k(x,\text{d}y)\mu_{1}(\text{d}x)$. Equivalently, $\mu$ decomposes into the conditional laws of the second coordinate given the first. This is typically guaranteed on standard Borel spaces (in particular, Polish spaces with their Borel $\sigma$-algebras).
@@
$P_{t}f(x)\coloneqq\mathbb{E}\left[f(X_{t})\mid X_{0}=x\right]=\int_{E}f(y)p_{t}(x,\text{d}y),\quad f\in B(E)\text{, }x\in E,$
where $B(E)$ is the space of bounded measurable functionals on $E$, and $p_{t}$ is the probability kernel describing the evolution of the Markov process. In words, operator $P_{t}$ applied on $f\in B(E)$ gives us the expected value of $f(X_{t})$, “weighted by” the transition probability (from the starting value $X_{0}=x$ to $X_{t}$) of our Markov process. We focus on this type of Markov semigroups because studying their (ergodic) behavior informs us about the (ergodic) behavior of the associated Markov processes.
Throughout the paper, we will assume that the semigroups we consider always have an invariant probability measure $\mu$, so that the elements of $\mathbf{P}=(P_{t})_{t\geq 0}$ are bounded operators in $\mathbb{L}^{p}(\mu)$, $p\in[1,\infty)$. In fact, most semigroups of interest have an invariant measure (Bakry et al., 2014, Section 1.2.1). A way to derive $\mu$ as a weak limit of a “reasonable” initial probability measure $\mu_{0}$ is inspected in Bakry et al. (2014, Page 10). For a discussion on this matter, we refer the interested reader to Remark 7 in Appendix C.1.
@@
Definition 2 (Imprecise Markov Semigroup (IMSG)).
An Imprecise Markov Semigroup is a nonempty collection $\mathbf{Q}$ of Markov semigroups (MSGs) associated with Markov processes. That is, $\mathbf{Q}\subset({B(E)}^{\mathbb{R}_{+}})^{B(E)}$ such that
@@
In Sections 3 and 4, we will work with compact IMSGs. To do so, we need to introduce a notion of compactness in the space $({B(E)}^{\mathbb{R}_{+}})^{B(E)}$. Such a notion hinges on the possibility of comparing the elements of $\mathbf{Q}$, when they are evaluated at some functional $\tilde{f}\in B(E)$.
First, suppose we are able to define a total preorder $\preceq^{\text{tot}}_{\tilde{f}}$ so that
@@
This relation is always reflexive and transitive on $\mathbf{Q}$; It need not be antisymmetric, since distinct semigroups may coincide on $\tilde{f}$. We therefore introduce the equivalence relation
$\mathbf{P}\sim_{\tilde{f}}\mathbf{P}^{\prime}\quad\Longleftrightarrow\quad P_{t}\tilde{f}=P_{t}^{\prime}\tilde{f}\text{ for all }t\geq 0,$
and work on the quotient set $\mathbf{Q}_{\tilde{f}}\coloneqq\mathbf{Q}/\sim_{\tilde{f}}$. The preorder $\preceq^{\mathrm{tot}}_{\tilde{f}}$ induces a genuine total order on $\mathbf{Q}_{\tilde{f}}$ by
$\mathbf{P}]\preceq^{\mathrm{tot}}_{\tilde{f}}[\mathbf{P}^{\prime}]\quad\Longleftrightarrow\quad\mathbf{P}\preceq^{\mathrm{tot}}_{\tilde{f}}\mathbf{P}^{\prime},$
@@
which is well-defined and antisymmetric by construction. Of course, here the square brackets denote equivalence classes. We have that, for all $\mathbf{P}],[\mathbf{P^{\prime}}]\in\mathbf{Q}_{\tilde{f}}$, either $\mathbf{P}]\preceq^{\text{tot}}_{\tilde{f}}[\mathbf{P^{\prime}}]$ or $\mathbf{P^{\prime}}]\preceq^{\text{tot}}_{\tilde{f}}[\mathbf{P}]$. In addition, the notions of maximal and greatest elements of $(\mathbf{Q}_{\tilde{f}},\preceq^{\text{tot}}_{\tilde{f}})$ coincide, and correspond to the classical notion of maximum (with respect to the total order $\preceq^{\text{tot}}_{\tilde{f}}$). Similarly, minimal and least element notions coincide, and correspond to the classical notion of minimum.
@@
Now, let $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$ be the order topology on $\mathbf{Q}_{\tilde{f}}$. That is (provided $|\mathbf{Q}_{\tilde{f}}|\geq 2$), $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$ is the topology whose base is given by the open rays $\{[\mathbf{P}]:[\mathbf{P}_{1}]\prec^{\text{tot}}_{\tilde{f}}[\mathbf{P}]\}$ and $\{[\mathbf{P}]:[\mathbf{P}]\prec^{\text{tot}}_{\tilde{f}}[\mathbf{P}_{2}]\}$, where $\prec^{\text{tot}}_{\tilde{f}}$ is the strict counterpart of $\preceq^{\text{tot}}_{\tilde{f}}$, for all $\mathbf{P}_{1}]$, $\mathbf{P}_{2}]$ in $\mathbf{Q}_{\tilde{f}}$, and by the open intervals $([\mathbf{P}_{1}],[\mathbf{P}_{2}])=\{[\mathbf{P}]:[\mathbf{P}_{1}]\prec^{\text{tot}}_{\tilde{f}}[\mathbf{P}]\prec^{\text{tot}}_{\tilde{f}}[\mathbf{P}_{2}]\}$. This means that the $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$-open sets in $\mathbf{Q}_{\tilde{f}}$ are the sets that are a union of (possibly infinitely many) such open intervals and rays.
@@
Assume that $\mathbf{Q}_{\tilde{f}}$ is $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$-compact. This implies that $\mathbf{Q}_{\tilde{f}}$ contains the minimum and the maximum (with respect to the total order $\preceq^{\text{tot}}_{\tilde{f}}$), which we denote by $\underline{[\mathbf{P}]}_{\tilde{f}}$ and $\overline{[\mathbf{P}]}_{\tilde{f}}$, respectively. We fix arbitrary representatives $\underline{\mathbf{P}}_{\tilde{f}}\in\underline{[\mathbf{P}]}_{\tilde{f}}$ and $\overline{\mathbf{P}}_{\tilde{f}}\in\overline{[\mathbf{P}]}_{\tilde{f}}$. In turn, since $\underline{\mathbf{P}}_{\tilde{f}},\overline{\mathbf{P}}_{\tilde{f}}\in\mathbf{Q}$, they both are (precise) Markov semigroups associated with Markov processes by (1).
@@
Suppose now that the induced relation is not total. More formally, the preorder $\preceq^{\text{part}}_{\tilde{f}}$ on $\mathbf{Q}$ is defined similarly to before by
$\mathbf{P}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime}}\iff P_{t}\tilde{f}\leq P^{\prime}_{t}\tilde{f}\text{, }\forall t\geq 0.$
Consider its induced relation (a genuine partial order) on the quotient $\mathbf{Q}_{\tilde{f}}$. Contrary to before, this time there might exist $\mathbf{P}],[\mathbf{P^{\prime}}]\in\mathbf{Q}_{\tilde{f}}$ such that $\mathbf{P}]\not\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P^{\prime}}]$ and $\mathbf{P}]\not\succeq^{\text{part}}_{\tilde{f}}[\mathbf{P^{\prime}}]$. A remark is in order.
Remark 1 (Antisymmetry and Identifiability).
Fix $\tilde{f}\in B(E)$ and consider the preorder on $\mathbf{Q}$ defined by